OpenStudy (anonymous):

Simplify quantity x plus 3 over 9 all over quantity x plus 12 over 6

OpenStudy (btaylor):

is this the right equation? \[\frac{\frac{x+3}{9}}{\frac{x+12}{6}}\] That is the same as \[\frac{x+3}{9} \times \frac{6}{x+12}\] See what you can do with that.